OpenStudy (anonymous):

I need help with a differential equation (1+y^2)(e^(2x) dx - e^y dy)-(1+y)dy = 0

OpenStudy (anonymous):

\[(1+y^2)(e ^{2x}dx - e^y dy) - (1 +y)dy=0\]

OpenStudy (rational):

that looks separable

OpenStudy (janu16):

Are thete any answer choices?

OpenStudy (anonymous):

\[e ^{2x}dx - e^y dy = \frac{ dy + y dy }{ 1+y^2 }\]

OpenStudy (anonymous):

This is something new we are seeing in class and I'm unsure as to how I should approach it.

OpenStudy (rational):

collect terms attached to `dx` on one side and terms attached to `dy` on other side

OpenStudy (janu16):

Is that thr equation or answer choice?

OpenStudy (anonymous):

no answer choice

OpenStudy (rational):

get it into this form : (stuff) `dy` = (stuff) `dx`

OpenStudy (anonymous):

\[e ^{2x}dx = \frac{ dy+y dy }{ 1+y^2 }+ e^y dy\]

OpenStudy (rational):

all `dy` s must be on one side

OpenStudy (rational):

all `dx` must be on other side

OpenStudy (anonymous):

I have all dx on the left side and dy on the right side

OpenStudy (rational):

Ahh nice :) simply integrate now

OpenStudy (rational):

\[\int e ^{2x}dx = \int \frac{ (1+y) dy }{ 1+y^2 }+\int e^y dy\]

OpenStudy (anonymous):

oo I see what you are doing.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }e^{2x}=\tan^{-1}(y)+\frac{ 1 }{ 2 } \ln|1+y^2|+e^y \]

OpenStudy (anonymous):

@rational just a little question, is this all to it?meaning that this falls as a "separable category" where it ends as a general solution such as x(x) + y (y) = C?